# What Do Latent Action Models Actually Learn?

### v1

@@ 4.1 Analysis 1: Linear LAM is PCA
We discuss the insight that this connection provides, followed by an analysis of several important cases covering specific settings of controllable changes and exogenous noise
@@
This proposition is equivalent to the Eckart–Young–Mirsky theorem (Eckart and Young, 1936) whose proof can be found in Chapter 2.4 of Golub and Van Loan (2013).
Over-parameterization issue. In fact, linear LAM is over-parameterized, resulting in the solution of $(A,B,C,D)$ for loss minimization being not unique. For example, the latent $\boldsymbol{z}=C\boldsymbol{o}+D\boldsymbol{o}^{\prime}$ is not necessarily confined to capturing information solely from $\boldsymbol{q}+\boldsymbol{\epsilon}$. Actually, it can be an arbitrary linear combination of the information from both $\boldsymbol{q}+\boldsymbol{\epsilon}$ and $\boldsymbol{o}$ since $A$ will later compensate to attenuate $\boldsymbol{o}$’s information in $\boldsymbol{z}$. To be more specific, there exists a family of solutions $B(C\boldsymbol{o}+D\boldsymbol{o}^{\prime})=(\boldsymbol{q}+\boldsymbol{\epsilon})+\alpha\boldsymbol{o}$ and $A=(1-\alpha)I$ for any $\alpha$ such that $\hat{\boldsymbol{o}}^{\prime}=B(C\boldsymbol{o}+D\boldsymbol{o}^{\prime})+A\boldsymbol{o}=(\boldsymbol{q}+\boldsymbol{\epsilon})+(1-\alpha)\boldsymbol{o}+\alpha\boldsymbol{o}=\boldsymbol{o}^{\prime}$.
We will revisit this issue in Section 4.3 and use data augmentation to handle this over-parameterization issue. For the ease of analysis before adopting this technique, we predict $(\hat{\boldsymbol{q}},\hat{\boldsymbol{\epsilon}},\hat{\boldsymbol{o}})$ from a surrogate latent $\tilde{\boldsymbol{z}}:=B^{-1}(\hat{\boldsymbol{o}}^{\prime}-\boldsymbol{o})$ when calculating LLO to get around this issue. This surrogate latent is the same as the original one when $A=I$, which is the case when data augmentation is adopted. Hence, we have $C=-D$ which indicates that the semantic meaning of the latent $\boldsymbol{z}=C\boldsymbol{o}+D\boldsymbol{o}^{\prime}$ remains the same (no movement) when $\boldsymbol{o}^{\prime}=\boldsymbol{o}$ across different observations.
Figure 3: LLO (Linear LAM objective) defined (6) measured when 1) $\boldsymbol{\epsilon}=0$ (left), 2) $\boldsymbol{\epsilon}$ is i.i.d. noise (middle), and 3) $\boldsymbol{\epsilon}$ contains the effect of other agents (right). Action MSE, noise MSE, and observation MSE are the three terms in (6) respectively. We set real action dimension $d_{a}$ and exogenous action dimension $d_{b}$ to $8$ in this experiment unless otherwise stated, and ensure $\boldsymbol{q}$ has unit variance.
Case 1: When $\boldsymbol{\epsilon}=0$. Without the interplay of the exogenous noise $\boldsymbol{\epsilon}$, linear LAM can capture the true action $\boldsymbol{a}$ in the latent $\boldsymbol{z}$ even if it does not have access to $\boldsymbol{a}$, when the bottleneck is set properly.
Considering that $\boldsymbol{\epsilon}=0$ and $\boldsymbol{q}=X\boldsymbol{a}$, the covariance matrix of $\Sigma_{\boldsymbol{q}+\boldsymbol{\epsilon}}=\Sigma_{\boldsymbol{q}}=\mathbb{E}[X\boldsymbol{a}\boldsymbol{a}^{T}X]$ only has $d_{a}$ non-zero eigenvalues. Following Proposition 4.2, we can make the following conclusions.
@@
Case 2: When $\boldsymbol{\epsilon}$ is i.i.d. noise. We start analyze the noise from the i.i.d. (independent and identically distributed) assumption, which is an approximation for many real noises such as the noise from the sensors or the image encoders. Now we assume $\boldsymbol{\epsilon}$ is i.i.d. noise with zero-mean $\mathbb{E}[\boldsymbol{\epsilon}]=0$ and covariance $\mathbb{E}[\boldsymbol{\epsilon}\boldsymbol{\epsilon}^{T}]=\sigma_{iid}^{2}I$. Considering that $\boldsymbol{q}$ and $\boldsymbol{\epsilon}$ are independent, the covariance matrix $\Sigma_{\boldsymbol{q}+\boldsymbol{\epsilon}}$ can be eigenvalue decomposed as $\Sigma_{\boldsymbol{q}+\boldsymbol{\epsilon}}=U\Gamma_{0}U^{T}=U_{0}(\Lambda_{0}+\sigma_{iid}^{2}I)U_{0}^{T}$ where $U_{0}$ and $\Lambda_{0}$ are the eigenvectors and eigenvalues of $\Sigma_{\boldsymbol{q}}$ respectively (i.e., when $\boldsymbol{\epsilon}=0$). Combining the results in Proposition 4.2, we conclude that, when there is i.i.d. noise, 1) the FDM parameter $B$ (and therefore the semantics of the latent since $B$ interpret the latent) remains the same, and 2) the loss increases since the eigenvalues increase. This conclusion is consistent with the conclusion on the robustness of PCA (Anderson, 1963; Johnstone, 2001).
@@
We present a simulation results on how LLO changes across different levels of the i.i.d. Gausion noise in Figure 3 (middle). We observe that linear LAM is robust to i.i.d. noise - even when the noise level $\sigma_{iid}=0.5$ (when the noise intensity is half that of the signal), linear LAM still succeeds with negligible gap to the optimal case.
@@
Case 3: When $\boldsymbol{\epsilon}$ contains the effect of other agents. In many practical datasets (such as Ego4d (Grauman et al., 2022)), the change between two observations is the effect of not only the controllable action but also other agents. For example, there may be another person walking in the background or camera shaking.
@@
Here, we assume $\boldsymbol{\epsilon}=Y\mathbf{b}$ where $Y\in\mathbb{R}^{d_{o}\times d_{b}}$ is the exogenous effect matrix and $\mathbf{b}$ represents the action taken by other agents. In this case, we can conduct an analysis similar to that of Case 1, and conclude that linear LAM will learn to capture the effects with top variances, no matter they result from the controllable action or other agents. Moreover, when the columns in $Y$ are not orthogonal to that of $X$, the FDM parameter $B$ will be impacted by the exogenous noise.
@@
We show the simulation results in Figure 3 (right) to illustrate which part of information (the controllable $\boldsymbol{q}$ or the noise $\boldsymbol{\epsilon}$) enters the latent when the latent dimension $d_{z}$ increases gradually. We can observe that: 1) When the variance of exogenous noise is smaller than $\boldsymbol{q}$ (i.e., $\sigma_{exo}=0.5$), linear LAM learns to fit the controllable part first, and $d_{z}=d_{a}$ is still the optimal configuration. 2) When the variance of noise exceeds that the signal $\boldsymbol{q}$ (i.e., $\sigma_{exo}=2.0$), linear LAM will first fit the noise, and the latent fails to exclude the information of the noise no matter how we set $d_{z}$. In this case, one possible solution is to clean training data to remove significant noises (such as shaking clips).
@@ 4.2 Analysis 2: Effects of Data Collection Policy
Till now, we consider the cases where both the controllable changes and the exogenous noise are independent of the observation, i.e., $\mathbb{E}[\boldsymbol{o}(\boldsymbol{q}+\boldsymbol{\epsilon})^{T}]=\mathbf{0}$. Unfortunately, this is usually not true for practical cases where LAM is trained using data collected by expert policies, e.g., many robot datasets are collected via teleoperation by humans. Within such datasets, the observation is correlated with the action $\boldsymbol{a}$, and thus the controllable change $\boldsymbol{q}$, resulting in $\mathbb{E}[\boldsymbol{o}(\boldsymbol{q}+\boldsymbol{\epsilon})^{T}]\neq\mathbf{0}$. Therefore, we dive into this scenario to see what is the effect of the data collection policy on linear LAM in this part.
Specifically, we assume the data is collected by a policy that generate action via $\boldsymbol{a}=\Pi_{d}\boldsymbol{o}+\boldsymbol{\pi}_{s}$ where $\Pi_{d}\in\mathbb{R}^{d_{a}\times d_{o}}$ represents the deterministic part of the policy and $\mathbf{\pi}_{s}\in\mathbb{R}^{d_{a}}$ represents the stochastic part of the policy. We assume that $\boldsymbol{o}$, $\boldsymbol{\epsilon}$, and $\boldsymbol{\pi}_{s}$ are uncorrelated.
In this case, since $\mathbb{R}[\boldsymbol{o}\boldsymbol{q}^{T}]\neq 0$, the cross term in the expansion of loss (7) remains. Letting $\Sigma_{\boldsymbol{o}}:=\mathbb{E}[\boldsymbol{o}\boldsymbol{o}^{T}]$ to be the covariance matrix of the observation, we can further solve for $A$ by setting the partial derivative of the loss function w.r.t. $A$ to zero, obtaining
@@
This differs from the previous result for the random policy where $A=I-(BC+BD)$ by the last term $\Sigma_{\boldsymbol{o}}\Pi_{d}^{T}X^{T}(BD-I)^{T}\Sigma_{\boldsymbol{o}}^{-1}$. The intuitive interpretation for this term is that $A$ can capture the changes caused by the deterministic part of the policy if these changes are not captured by the latent (noting that $BD-I$ can project the vectors onto the orthogonal complement of the column space of $B$). In other words, the latent may not capture the effect of each real action. Instead, the FDM would capture the effect of the deterministic part of the policy, whereas the latent would capture the remaining stochastic part.
In our simulation, we control the randomness of the policy using a coefficient $\chi$, and let $\boldsymbol{a}=\chi\Pi_{d}\boldsymbol{o}+(1-\chi)\boldsymbol{\pi}_{s}$ where $\Pi_{d}$ is a random projection matrix and $\pi_{s}$ is a standard Gaussian vector. The simulation results in Figure 4 (left) indicate that the more deterministic the data collection policy is, the less information about $\boldsymbol{a}$ is captured by $\boldsymbol{z}$. Therefore, the analysis in this section suggest us to use the data collected by the policies with higher randomness whenever possible. This may also indicate that, for LAM training, the trajectories collected by expert policies may not be enough, and more exploratory trajectories are needed.
Figure 4: LLO (Linear LAM objective) defined (6) measured when 1) data is collected by policies of different deterministic levels (left), 2) linear LAM is trained w/ and w/o data augmentation (middle), and 3) linear LAM is trained w/ and w/o action prediction (right). We set $d_{a}=8$ in the experiments.
@@ 4.3 Analysis 3: Improvements via Data Augmentation
This section analyzes the augmentation scheme used in Chen et al. (2024b), which will turn out to resolve a key issue encountered in linear LAM (see the discussion on over-parameterization in Section 4.1), where information about the observation $\boldsymbol{o}$ enters into the latent $\boldsymbol{z}$.
The IGOR model (Chen et al., 2024b) applies one shared random crop to both inputs of the IDM, and a second shared random crop applied to the FDM and reconstruction target. The intuition is that “by using different croppings [for IDM and FDM], the model is encouraged to learn a more semantically invariant latent action”. Our subsequent analysis shows this intuition is provably correct in linear LAM, resulting in new terms to the loss that encourage removal of the latent of any semantic information about the observation. Sun et al. (2024) also apply a similar scheme, with a single action preserving crop applied to the IDM which makes it harder “to copy the appearance information directly from the [IDM]”.
Data augmentation in linear LAM. We extend our linear LAM setup to include a data augmentation operator $\operatorname{Aug}_{i}[\boldsymbol{o}]\coloneqq\boldsymbol{o}+\boldsymbol{\kappa}_{i}$ that takes an observation as input and adding some random vector $\boldsymbol{\kappa}\in\mathbb{R}^{d}_{o}$. $\operatorname{Aug}_{i}[\cdot]$ applies the same $i$-th operator to different variables, say $\operatorname{Aug}_{1}[\boldsymbol{o}]$ and $\operatorname{Aug}_{1}[\boldsymbol{o}^{\prime}]$ will apply a consistent random variable $\boldsymbol{\kappa}_{1}$ to each observation.
@@
We provide the proof in Appendix C.2. What does it mean for these terms to encourage $A=I$ and $C+D=0$? If, $A$ is the identity, all observation information flows directly through this matrix, and $\boldsymbol{z}$ need not carry any additional information about $\boldsymbol{o}$, under the additive assumption on the dynamics. With similar effect, setting $C+D=0$ immediately cancels out $\boldsymbol{o}$ information in $\boldsymbol{z}$.
@@
This condition also hints semantic consistency – the semantic meaning of the latent the frame does not change from $\boldsymbol{o}$ to $\boldsymbol{o}^{\prime}$ will be consistent across different observations. Hence, this augmentation scheme is a mechanism to remove information about $\boldsymbol{o}$ from $\boldsymbol{z}$, explicitly minimizing $\mathcal{I}(\boldsymbol{z};\boldsymbol{o})$ in the original objective (5). In this way, data augmentation results in latents with better semantics.
Simulation. We show the effect of data augmentation under different noise level in Figure 4 (middle) The results indicate the by data augmentation can improve the semantics of the latent $\boldsymbol{z}$ by addressing the over-parameterization issue. Note that, starting here, the simulation calculate LLO based on the true latent $\boldsymbol{z}$ instead of the pseudo-latent. In this way, even when there is no noise $\sigma_{iid}=0$, the learned latent $\boldsymbol{z}$ cannot perfectly predict the real action $\boldsymbol{a}$ without involving other information. However, adding in the augmentation vector with 10% variance (compared with that of observation $\boldsymbol{o}$), can greatly improve the learning of linear LAM.
Designing data augmentation. In our linear setting, a natural design for data augmentation is to have $\boldsymbol{\kappa}$ i.i.d. across its elements. This design is based on our knowledge that the semantic meaning for the frame change should be invariant to this data augmentation, i.e., $\boldsymbol{o}^{\prime}-\boldsymbol{o}=(\boldsymbol{o}^{\prime}+\boldsymbol{\kappa})-(\boldsymbol{o}+\boldsymbol{\kappa})$, since we know the dynamics is additive. For real images, Chen et al. (Chen et al., 2024b) adopts random crops, reflecting the prior that the semantic meaning about frame change is invariant to this way of augmentation. Further, the variance of different augmentations determines how important this term is – a larger augmentation variance enforces this constraint more strictly. This indicates that we may design a weighted mix of data augmentation that best capture the invariance property to improve the semantics of the learned latent.

### v3

@@ 4.1 Analysis 1: Linear LAM is PCA
We discuss the insight that this connection provides, followed by an analysis of several important cases covering specific settings of controllable changes and exogenous noise. Figure 5 gives intuitive examples of these noise types. Figure 4 summarizes the technical results.
@@
This proposition is equivalent to the Eckart–Young–Mirsky theorem (eckart1936pca) whose proof can be found in Chapter 2.4 of golub2013matrix.
Over-parameterization issue. We show that linear LAM is over-parameterized, with multiple solutions of $(A,B,C,D)$ able to minimize the reconstruction loss objective. Specifically, the latent $\boldsymbol{z}=C\boldsymbol{o}+D\boldsymbol{o}^{\prime}$, in addition to capturing information about $\boldsymbol{q}$ and $\boldsymbol{\epsilon}$, may further contain information about $\boldsymbol{o}$. There is no detrimental effect provided the $A$ matrix compensates to ‘knock out’ $\boldsymbol{o}$’s information in $\boldsymbol{z}$. Concretely, there exists a family of solutions $B(C\boldsymbol{o}+D\boldsymbol{o}^{\prime})=(\boldsymbol{q}+\boldsymbol{\epsilon})+\alpha\boldsymbol{o}$ and $A=(1-\alpha)I$ for any $\alpha\in\mathbb{R}$ such that $\hat{\boldsymbol{o}}^{\prime}=B(C\boldsymbol{o}+D\boldsymbol{o}^{\prime})+A\boldsymbol{o}=(\boldsymbol{q}+\boldsymbol{\epsilon})+(1-\alpha)\boldsymbol{o}+\alpha\boldsymbol{o}=\boldsymbol{o}^{\prime}$.
[…]
Figure 3: LLO (Linear LAM objective (6), higher better) measured in three noise settings. (Left) $\boldsymbol{\epsilon}=0$. (Middle) $\boldsymbol{\epsilon}$ is i.i.d. noise. (Right) $\boldsymbol{\epsilon}$ contains the effect of other agents. Action MSE, noise MSE, and observation MSE are the three terms in (6). We set real action dimension $d_{a}=8$ and exogenous action dimension $d_{b}=8$ unless otherwise stated, and ensure $\boldsymbol{q}$ has unit variance.
Case 1: $\boldsymbol{\epsilon}=0$. In the absence of exogenous noise $\boldsymbol{\epsilon}$, linear LAM does capture the true action $\boldsymbol{a}$ in the latent $\boldsymbol{z}$, when the bottleneck dimension is set equal or larger than the action dimension.
When $\boldsymbol{\epsilon}=0$ and $\boldsymbol{q}=X\boldsymbol{a}$, the covariance matrix of $\Sigma_{\boldsymbol{q}+\boldsymbol{\epsilon}}=\Sigma_{\boldsymbol{q}}=\mathbb{E}[X\boldsymbol{a}\boldsymbol{a}^{T}X]$ only has $d_{a}$ non-zero eigenvalues. Following Proposition 4.2, we can make the following conclusions.
@@
Case 2: $\boldsymbol{\epsilon}$ is i.i.d. noise. We consider i.i.d. noise (independent and identically distributed), which may be a realistic assumption in the case of sensors or image encoders. We assume $\boldsymbol{\epsilon}$ is i.i.d. with zero-mean $\mathbb{E}[\boldsymbol{\epsilon}]=0$ and isotropic covariance $\mathbb{E}[\boldsymbol{\epsilon}\boldsymbol{\epsilon}^{T}]=\sigma_{iid}^{2}I$. Considering that $\boldsymbol{q}$ and $\boldsymbol{\epsilon}$ are independent, the covariance matrix $\Sigma_{\boldsymbol{q}+\boldsymbol{\epsilon}}$ can be eigenvalue decomposed as $\Sigma_{\boldsymbol{q}+\boldsymbol{\epsilon}}=U\Gamma_{0}U^{T}=U_{0}(\Lambda_{0}+\sigma_{iid}^{2}I)U_{0}^{T}$ where $U_{0}$ and $\Lambda_{0}$ are the eigenvectors and eigenvalues of $\Sigma_{\boldsymbol{q}}$ respectively (i.e., when $\boldsymbol{\epsilon}=0$). Combining the results in Proposition 4.2, we conclude that, when there is i.i.d. noise, 1) the FDM parameter $B$ (and therefore the semantics of the latent since $B$ interpret the latent) remains the same, and 2) the loss increases since the eigenvalues increase. This conclusion is consistent with the conclusion on the robustness of PCA (anderson1963asymptotic; johnstone2001distribution).
@@
Figure 3 (middle) shows how LLO changes for Gaussian noise of differing variance. We observe that linear LAM is robust to i.i.d. noise up to around $\sigma_{iid}=0.5$ (when the noise intensity is half that of the signal), linear LAM still succeeds with negligible gap to the optimal case.
@@
Case 3: $\boldsymbol{\epsilon}$ contains the effect of other agents. In many real-world datasets (such as Ego4d (grauman2022ego4d)), the change between two observations may not only be caused by the control action of the ego-agent (e.g. joints of a robot arm), but additionally can be effected by exogenous noise, such as ‘other agents’ (e.g. a person walking in the background, camera shake) – see Figure 5. Compared with i.i.d. noise, this noise is structured and would be predictable if the other agent’s control actions were available.
@@
Here, we assume $\boldsymbol{\epsilon}=Y\mathbf{b}$ where $Y\in\mathbb{R}^{d_{o}\times d_{b}}$ is the exogenous effect matrix and $\mathbf{b}$ represents the action taken by other agents. In this case, analysis similar to that of Case 1 concludes that linear LAM will learn to capture effects with largest variances, no matter if they result from the controllable action or other agents. Moreover, when the columns in $Y$ are not orthogonal to that of $X$, the FDM parameter $B$ will be impacted by the exogenous noise.
@@
Simulation results in Figure 3 (right) illustrate which part of information (the controllable $\boldsymbol{q}$ or the noise $\boldsymbol{\epsilon}$) enters the latent when the latent dimension $d_{z}$ is increased. We observe that: 1) When the variance of exogenous noise is smaller than $\boldsymbol{q}$ (i.e., $\sigma_{q}=1>\sigma_{exo}=0.5$), linear LAM learns to fit the controllable part first, and $d_{z}=d_{a}$ is still the optimal configuration. 2) When the variance of noise exceeds that the signal $\boldsymbol{q}$ (i.e., $\sigma_{q}=1<\sigma_{exo}=2.0$), linear LAM will first fit the noise, and the latent fails to exclude the information of the noise no matter how we set $d_{z}$. To alleviate this issue, an obvious solution in practice is to preprocess training data to reduce significant noise components (such as stabilizing camera shake).
Figure 4: Visualizaton of the technical results in Section 4.1, when correlation between $\boldsymbol{o},\boldsymbol{q}$ and $\mathbf{\epsilon}$ are assumed zero, variance decomposes additively. As with PCA, linear LAM’s latent indiscriminately captures the $d_{z}$ dimensions with largest variance (i.e., the top dimensions in the figure), regardless of the source of this variance.
Figure 5: Intuitive examples of the noise cases defined in Section 4.1. Images are taken from the RT-1 (the first two examples) and Ego4d (the last example) datasets. Differences in subsequent frames in the RT-1 dataset are dominated by the controllable actions, while the camera shake in Ego4d leads to large visual changes not related to the actions of interest (here hand manipulation).
@@ 4.2 Analysis 2: Effects of Data Collection Policy
Section 4.1 considered the case where both the controllable changes and the exogenous noise are independent of the observation, i.e., $\mathbb{E}[\boldsymbol{o}(\boldsymbol{q}+\boldsymbol{\epsilon})^{T}]=\mathbf{0}$. However, practical LAMs are usually trained on data created by expert policies, e.g., robotic datasets often consist of human teleoperation. Within such datasets, the observation is correlated with the action $\boldsymbol{a}$, and thus the controllable change $\boldsymbol{q}$, resulting in $\mathbb{E}[\boldsymbol{o}(\boldsymbol{q}+\boldsymbol{\epsilon})^{T}]\neq\mathbf{0}$. This section analyzes this case in linear LAM.
Specifically, we assume the data generating policy acts via $\boldsymbol{a}=\Pi_{d}\boldsymbol{o}+\boldsymbol{\pi}_{s}$ where $\Pi_{d}\in\mathbb{R}^{d_{a}\times d_{o}}$ represents the deterministic part of the policy and $\mathbf{\pi}_{s}\in\mathbb{R}^{d_{a}}$ represents the stochastic part of the policy. We assume that $\boldsymbol{o}$, $\boldsymbol{\epsilon}$, and $\boldsymbol{\pi}_{s}$ are uncorrelated, but there will be correlation between $\boldsymbol{o}$ and $\boldsymbol{a}$.
Return to the expanded loss objective (Appendix C.1 (18)),
$\mathcal{L}=$ $\lVert(A+BC+BD-I)\boldsymbol{o}\rVert_{2}^{2}+\lVert(BD-I)(\boldsymbol{q}+\boldsymbol{\epsilon})\rVert_{2}^{2}$
$\qquad\qquad\qquad\qquad\qquad\quad+2\operatorname{Tr}\left((A+BC+BD-I)\boldsymbol{o}(\boldsymbol{q}+\boldsymbol{\epsilon})^{T}(BD-I)\right).$ (8)
Since $\mathbb{R}[\boldsymbol{o}\boldsymbol{q}^{T}]\neq 0$, the cross term in the expansion of this loss cannot be ignored, and no longer reduces to PCA. Letting $\Sigma_{\boldsymbol{o}}:=\mathbb{E}[\boldsymbol{o}\boldsymbol{o}^{T}]$ be the covariance matrix of the observation, we can further solve for $A$ by setting the partial derivative of the loss function w.r.t. $A$ to zero, obtaining
@@
This differs from the previous result for the random policy where A=I(BC+BD)A=I-(BC+BD) by the last term $\Sigma_{\boldsymbol{o}}\Pi_{d}^{T}X^{T}(BD-I)^{T}\Sigma_{\boldsymbol{o}}^{-1}$. The intuitive interpretation for this term is that $A$ will capture changes caused by the deterministic part of the policy (noting that BDIBD-I can project the vectors onto the orthogonal complement of the column space of $B$). This is problematic for LAM, since the FDM can implicitly absorb deterministic parts of the policy, leaving capacity in the latent available to absorb exogenous noise. Intuitively, if action $\boldsymbol{a}^{*}$ is always selected for observation $\boldsymbol{o}^{*}$ , this always leads to $\boldsymbol{o}^{\prime*}$. The FDM can simply learn that the mapping $\boldsymbol{o}^{*}\to\boldsymbol{o}^{\prime*}$, without the need for a latent $\boldsymbol{z}$.
In our simulation, we control the randomness of the policy using a coefficient $\chi$, and let $\boldsymbol{a}=\chi\Pi_{d}\boldsymbol{o}+(1-\chi)\boldsymbol{\pi}_{s}$ where $\Pi_{d}$ is a random projection matrix and $\boldsymbol{\pi_{s}}$ is a standard Gaussian vector. The simulation results in Figure 6 (left) indicate that the more deterministic the data collection policy is, the less information about $\boldsymbol{a}$ is captured by $\boldsymbol{z}$. Therefore, the analysis in this section suggests data collection policies with higher randomness are preferred for LAM training. Trajectories collected by deterministic expert policies may lead to poor quality LAMs.
Figure 6: LLO (6) in three settings. (Left) Data generated from policies ranging from fully random to fully deterministic. (Middle) Linear LAM trained w/ varying strength data augmentation. (Right) Linear LAM trained w/ and w/o action prediction. We set $d_{a}=8$ in the experiments.
@@ 4.3 Analysis 3: Improvements via Data Augmentation
This section analyzes the augmentation scheme used in chen2024igor, which will turn out to resolve a key issue encountered in linear LAM (see the discussion on over-parameterization in Section 4.1), where information about the observation $\boldsymbol{o}$ enters into the latent $\boldsymbol{z}$.
The IGOR model (chen2024igor) applies one shared random crop to both inputs of the IDM, and a second shared random crop applied to the FDM and reconstruction target. The intuition is that “by using different croppings [for IDM and FDM], the model is encouraged to learn a more semantically invariant latent action”. Our subsequent analysis shows this intuition is provably correct in linear LAM, resulting in new terms to the loss that encourage removal of the latent of any semantic information about the observation. sun2024deltadiffusion also apply a similar scheme, with a single action preserving crop applied to the IDM which makes it harder “to copy the appearance information directly from the [IDM]”.
Data augmentation in linear LAM. We extend our linear LAM setup to include a data augmentation operator $\operatorname{Aug}_{i}[\boldsymbol{o}]\coloneqq\boldsymbol{o}+\boldsymbol{\kappa}_{i}$ that takes an observation as input and adds some random vector $\boldsymbol{\kappa}\in\mathbb{R}^{d}_{o}$. $\operatorname{Aug}_{i}[\cdot]$ applies the same $i$-th operator to different variables, say $\operatorname{Aug}_{1}[\boldsymbol{o}]$ and $\operatorname{Aug}_{1}[\boldsymbol{o}^{\prime}]$ will apply a consistent random variable $\boldsymbol{\kappa}_{1}$ to each observation.
@@
See Appendix C.2 for proof. What is the effect of encouraging $A=I$ and $C+D=0$? If, $A$ is the identity, all observation information flows directly through this matrix, and $\boldsymbol{z}$ need not carry any additional information about $\boldsymbol{o}$ (given our additive transition function). Equivalently, setting $C+D=0$ cancels out $\boldsymbol{o}$ information in $\boldsymbol{z}$.
@@
This condition improves the semantic meaning of the latent action across different observations. For example, the zero latent $\mathbf{z=0}$ has the semantic meaning “the frame does not change from $\boldsymbol{o}$ to $\boldsymbol{o}^{\prime}$, which is consistent across different observations. Hence, this augmentation scheme is a mechanism to remove information about $\boldsymbol{o}$ from $\boldsymbol{z}$, explicitly minimizing $\mathcal{I}(\boldsymbol{z};\boldsymbol{o})$ in the original objective (5).
Simulation. We show the effect of data augmentation under different noise levels in Figure 6 (middle). Starting from here, the simulations calculate LLO based on the true latent $\boldsymbol{z}$ instead of the pseudo-latent. Due to this change, even when there is no noise $\sigma_{iid}=0$, the learned latent $\boldsymbol{z}$ cannot perfectly predict the real action $\boldsymbol{a}$ without other information. However, adding an augmentation vector with 0.1 magnitude variance (10% of the observation variance), greatly improves the latent learned by linear LAM, achieving close to the ideal LLO.
Designing data augmentation. In our linear setting, we have designed the data augmentation $\boldsymbol{\kappa}$ as additive and i.i.d. across its elements. This design is based on our knowledge that the semantic meaning for the frame change should be invariant to this data augmentation, i.e., $\boldsymbol{o}^{\prime}-\boldsymbol{o}=(\boldsymbol{o}^{\prime}+\boldsymbol{\kappa})-(\boldsymbol{o}+\boldsymbol{\kappa})$, since the dynamics are also additive. For real images, chen2024igor adopt random crops, reflecting a prior that the action is invariant to position (e.g., a robot closing its gripper should produce the same latent action regardless of its location in the image). Further, the variance of different augmentations determines how important this term is – a larger augmentation variance enforces this constraint more strictly. In summary, our analysis justifies the practice of designing data augmentations that capture a model designer’s domain knowledge about desired invariances to improve the semantics of the learned latent.